OpenStudy (anonymous):

Which expression is not equivalent to the other expressions? zero point five times four raised to the second power plus ten two point five times the quantity six minus two raised to the second power, end quantity two raised to the third power divided by ten plus four point two the quantity six plus three raised to the second power, end quantity, divided by three

OpenStudy (anonymous):

0.5×4^2+10 = 18 2.5×(6−2^2) = 5 2^3÷10+4.2 = 5 (6+3^2)÷3 = 5

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

thank you for the medal